The area of a right triangle s 44cm. find the lenght of the legs if one is 3 cm longer than the other
Answer by scott8148(6628)

(1/2)(L)(L + 3) = 44 ___ L^2 + 3L = 88 ___ L^2 + 3L - 88 = 0

(L - 8)(L + 11) = 0